### v3

# Whisper-Flamingo: Integrating Visual Features into Whisper for Audio-Visual Speech Recognition and Translation

Abstract
Audio-Visual Speech Recognition (AVSR) uses lip-based video to improve performance in noise. Since videos are harder to obtain than audio, the video training data of AVSR models is usually limited to a few thousand hours. In contrast, speech models such as Whisper are trained with hundreds of thousands of hours of data, and thus learn a better speech-to-text decoder. The huge training data difference motivates us to adapt Whisper to handle video inputs. Inspired by Flamingo which injects visual features into language models, we propose Whisper-Flamingo which integrates visual features into the Whisper speech recognition and translation model with gated cross attention. Our models achieve state-of-the-art ASR WER (0.68%) and AVSR WER (0.76%) on LRS3, and state-of-the-art ASR WER (1.3%) and AVSR WER (1.4%) on LRS2. Audio-visual Whisper-Flamingo outperforms audio-only Whisper on English speech recognition and En-X translation for 6 languages in noisy conditions. Moreover, Whisper-Flamingo is versatile and conducts all of these tasks using one set of parameters, while prior methods are trained separately on each language.
keywords:
audio-visual speech recognition, noise-robust
1 Introduction
In recent years, major improvements in Automatic Speech Recognition (ASR) performance have been achieved by models trained on large-scale data [1, 2], but performance still declines in noise [3]. To enhance performance in noise, Audio-Visual Speech Recognition (AVSR) uses lip-based video in addition to audio inputs. Self-Supervised Learning (SSL) methods such as AV-HuBERT [4] pre-train on large-scale datasets of unlabeled videos and fine-tune on a few hundred hours of labeled videos to perform noise-robust AVSR. However, due to the difficulty in collecting publicly accessible videos, these models are usually trained on only a few thousand hours of data.
To overcome the lack of video data, recent methods fine-tune audio-only models pre-trained on hundreds of thousands of hours of audio for audio-visual speech recognition [5, 6, 7]. The results show that such audio models combined with audio-visual fine-tuning on a few hundred hours of videos can approach the performance of video models pre-trained on thousands of hours of video [6]. However, these methods often train video models and text decoders from scratch on only a few hundred hours of data, which is suboptimal compared to training on large-scale data. Furthermore, only English data has been used.
In this work, we propose to integrate visual features from AV-HuBERT into Whisper [1], an audio-only model trained on 680k hours of speech with a strong multilingual decoder. Compared to the prior audio-visual adaptation methods, our video model and text decoder are pre-trained with large-scale data. This allows our method to perform well on audio-visual speech translation, a task not explored by prior methods [5, 6, 7].
How to fuse modalities effectively in multi-modal models is an ongoing research question. One recent work, Flamingo [8], fuses visual features into text-only language models using gated cross attention and fine-tuning on a paired text-image dataset. The gated cross attention layers are initialized to the identity function and learn to attend to the visual features during fine-tuning. These layers have been shown to generalize to different modality pairs; Audio Flamingo [9] recently applied them for text-audio reasoning. Inspired by this method, we propose Whisper-Flamingo which inserts gated cross attention layers into Whisper’s decoder and enables Whisper to use lip-based features for speech recognition.
On the English (En) LRS3 video dataset [10], our models achieve State-of-the-Art (SOTA) ASR WER (0.68%) and AVSR WER (0.76%). On LRS2 [11], our models achieve SOTA ASR WER (1.3%) and AVSR WER (1.4%). Our novel audio-visual Whisper-Flamingo significantly outperforms the audio-only Whisper baseline in noise. Moreover, Whisper-Flamingo achieves competitive noise-robust results compared to prior audio-visual models. Next, we demonstrate Whisper’s multilingual capabilities by extending Whisper-Flamingo for En-X translation on the MuAViC dataset [12]. Our model performs En transcription and En-X translation into 6 other languages, while the previous audio-visual SOTA requires fine-tuning on each language separately. Once again, Whisper-Flamingo significantly outperforms audio-only Whisper in noise for both En transcription and En-X translation. Code and models at https://github.com/roudimit/whisper-flamingo
Figure 1: Diagram of Whisper-Flamingo based on Whisper [1] and Flamingo [8]. We first fine-tune all of Whisper’s parameters using English audio for English transcription and En-X translation. To train Whisper-Flamingo, we freeze the audio model, add gated cross attention layers into Whisper’s decoder attending to visual features from AV-HuBERT, and train the model on audio-visual inputs.
2 Method
In this section, we review audio-visual fusion methods for AVSR, and then explain our method. Two common fusion methods are early and late fusion. In early fusion, both modalities are first separately processed by light-weight encoders and then combined with feature addition or concatenation and used as input to an audio-visual Transformer [13, 11]. Both SSL models [4, 14, 15] and fully-supervised models [16, 17, 18] use this design. In late fusion, audio and video are processed separately by Transformer encoders, and afterwards features are fused with an MLP. The audio-visual features are then passed to a linear-layer or Transformer decoder. This approach is common for fully-supervised models [19, 20, 21]. Both early and late fusion need identical audio and visual feature rates so that they can be fused at each time step; a common design is to match the video’s frame rate by downsampling the audio features to 25 Hz.
Table 1: Hyperparameter summary. We first train Whisper-Large FT (Fine-tune) with audio-only, then use it to initialize Whisper-Flamingo. A=audio, AV=audio-visual. †=per sample.
Whisper- Whisper- Whisper- Whisper-
Large FT Large FT Flamingo Flamingo
Test Modalities A A AV AV
En Recognition
En-X Translation
GPUs 1 4 1 4
Total Params. 1.55B 1.55B 2.5B 2.5B
AV-HuBERT Params. - - 325M 325M
Gated X-Attn Params. - - 630M 630M
Trainable Params. 1.55B 1.55B 631M 631M
Warmup Steps 1k 1k 5k 5k
Total Steps 90k 225k 20k 40k
Learning Rate $5\times 10^{-6}$ $5\times 10^{-6}$ $1\times 10^{-4}$ $1\times 10^{-4}$
Batch per GPU (s) 80 40 160 30
Max Length (s) † 10 10 15 15
Max Characters † 350 300 350 250
Most methods which adapt pre-trained audio-only models for AVSR through audio-visual fine-tuning use early fusion. FAVA [6] adapts BEST-RQ [22], an audio self-supervised model, through early fusion with a video model trained from scratch. Adaptive AV [7] prepends Whisper with an audio-visual Transformer to output a de-noised spectrogram, but does not use visual features directly in Whisper. However, we found gated cross attention with features from pre-trained AV-HuBERT worked better than early fusion. Note that separate research focuses on using visual features from images or instructional videos for AVSR, where the visuals provide context and are only loosely synchronized with the audio [23, 24].
We propose to adapt Whisper’s decoder with visual features from AV-HuBERT using gated cross attention, as shown in Figure 1. Each of Whisper’s decoder blocks consists of a self-attention layer, cross attention layer attending to the audio features, and a Multi-Layer Perceptron (MLP). Based on Flamingo [8], the gated cross attention layer is defined as follows, where $\mathbf{x}$ is the input to the decoder block, $\mathbf{v}$ are the visual features, $\mathrm{Attn}$ is multi-head cross attention, $\mathrm{LN}$ is Layernorm [25], and $\mathrm{FFW}$ is an MLP:
$\mathbf{x^{\prime}}$ $=\mathbf{x}+\mathrm{tanh}(\mathrm{\alpha_{xattn}})\times\mathrm{Attn}(\mathrm{LN}(\mathbf{x}),\mathbf{v})$ (1)
$\mathbf{y}$ $=\mathbf{x^{\prime}}+\mathrm{tanh}(\mathrm{\alpha_{mlp}})\times\mathrm{FFW}(\mathrm{LN}(\mathbf{x^{\prime}}))$ (2)
The learnable parameters $\mathrm{\alpha_{xattn}}$ and $\mathrm{\alpha_{mlp}}$ are initialized to 0 so that the layers initially function as the identity since $\tanh(0)=0$. Through audio-visual fine-tuning, the model adjusts the weights of $\mathrm{\alpha_{xattn}}$ and $\mathrm{\alpha_{mlp}}$ and learns to attend to the visual features. We insert the gated cross attention layers in Whisper’s decoder in the beginning of each block, before the self-attention layer. We tried to insert them in other orders within the decoder blocks, but the performance was slightly worse. The full analysis is shown in the Appendix, Section 7.4. Note that since the gated cross attention separately attends to the video features, the audio and video features can have different feature rates (for example, 50 Hz and 25 Hz).
Table 2: Fusion ablation with Whisper-Medium on LRS3. We report results on the original test set (Clean) and with babble noise injected at 0-SNR (Noisy). A=audio, AV=audio-visual.
Test Clean Noisy
[…]
Whisper, Zero-shot A 2.3 22.2
Whisper, Fine-tuned A 1.9 12.6
Whisper-Early-Fusion AV 1.7 10.0
Whisper-Late-Fusion AV 2.1 16.5
Whisper-Flamingo AV 1.5 7.0
Table 3: Results for English transcription on LRS3. We report results on the original test set (Clean) and with babble noise added at 0-SNR (Noisy). Hours of unlabeled & labeled audio-visual data used to train each model are shown. 433h=LRS3, 1,759h=LRS3+VoxCeleb2. Noise dataset= dataset used to make babble noise for testing. Noisy WER from methods using different noise datasets are not directly comparable. Note that u-HuBERT [14], AV-HuBERT [26], and CMA [27] use a different noise file than us. *=noisy results were not reported using 0-SNR.
[…]
Model Dataset Unlabeled Labeled Clean Noisy Clean Noisy
Audio-Visual methods trained only on transcribed videos
VIT 3D [17] NoiseX - 90k 1.6 6.1 1.6 2.9
LP Conformer [28] NoiseX - 100k - - 0.9 1.9
AutoAVSR [21] NoiseX - 1,902 / 3,448 1.0 / 1.0 * 1.0 / 0.9 *
Fast Conformer [29] NoiseX - 435 / 3,116 1.6 / 0.7 * 1.7 / 0.8 *
Audio-Visual SSL methods
AV2vec [15] MUSAN 433 433 2.7 19.5 2.5 6.7
AV-BEST-RQ [6] NoiseX 1,759 433 - - 2.1 6.8
BRAVEn [30] NoiseX 2,649 433 1.1 - - -
AV-HuBERT [26] LRS3† 1,759 433 1.6 15.7 1.4 5.8
u-HuBERT [14] LRS3† 1,759 433 1.4 19.3 1.3 4.6
CMA [27] LRS3† 1,759 433 - - 1.5 4.4
Audio-Only pre-train + audio-visual fine-tune methods
Adaptive AV [7] MUSAN 400 30 - - 2.3 16.3
FAVA-USM [6] NoiseX - 433 - - 1.3 6.2
Llama-AVSR [31] NoiseX 1,759 433 / 1,759 1.1 / 0.81 12.3 / - 0.95 / 0.77 4.2 / -
Our audio-only Whisper Zero-shot baseline
Whisper-Large, Zero-shot, Beam 1 LRS3 - - 2.3 23.3 - -
$\hookrightarrow$ w/ Beam search 15 LRS3 - - 2.1 20.8 - -
Our models trained without noise
Whisper-Large, Fine-tuned w/o noise, Beam 1 LRS3 - 433 / 1,759 1.0 / 0.68 21.6 / 23.6 - -
$\hookrightarrow$ w/ Beam search 15 LRS3 - 433 / 1,759 1.3 / 1.5 23.1 / 24.1 - -
Whisper-Flamingo, Fine-tuned w/o noise, Beam 1 LRS3 1,759 433 / 1,759 - - 1.1 / 0.76 13 / 14.9
$\hookrightarrow$ w/ Beam search 15 LRS3 1,759 433 / 1,759 - - 1.1 / 1.8 12.8 / 15.9
Our models trained with noise
Whisper-Large, Fine-tuned w/ noise, Beam 1 LRS3 - 433 / 1,759 1.1 / 0.85 12.4 / 11.9 - -
$\hookrightarrow$ w/ Beam search 15 LRS3 - 433 / 1,759 2.3 / 2.0 11.7 / 11.1 - -
Whisper-Flamingo, Fine-tuned w/ noise, Beam 1 LRS3 1,759 433 / 1,759 - - 1.0 / 0.86 7.2 / 6.1
$\hookrightarrow$ w/ Beam search 15 LRS3 1,759 433 / 1,759 - - 1.5 / 2.0 5.6 / 5.6
Training pipeline. Before adding gated cross attention, we first fine-tune all layers of the audio-only Whisper model to adapt it to the domain of interest (denoted as Whisper Fine-tuned). We also add noise during fine-tuning to increase the noise-robustness. We use the standard cross-entropy loss between the model’s predicted transcript and the ground-truth tokens. To train Whisper-Flamingo, we freeze the fine-tuned Whisper, insert the gated cross attention layers, and fine-tune the model with audio-visual inputs. The gated cross attention layers and a linear layer on top of the visual features are trained from scratch, while all other parameters are frozen. The new layers can therefore be seen as a (large) set of adaptors [32]: removing them results in the audio-only Whisper weights.
From English to Multilingual. Whisper was trained for multilingual transcription and X-En translation (multilingual audio to En text). We tried Whisper-Flamingo on multilingual speech recognition and X-En translation using the videos in the MuAViC dataset [12] but found several issues. Most languages in the dataset have less than a third of the hours of English data available, which makes training new layers from scratch difficult. Also, the multilingual videos are longer on average than the English videos. This causes increased GPU memory pressure and requires a reduced batch size, which also makes training difficult. Therefore we focused on En-X translation (English audio to multilingual text) and propose to handle multilingual recognition and translation in future work [33, 34, 35, 29, 36].
Prior research shows that Whisper can be prompted for En-X translation, but it requires language-specific logit filtering and the performance can still be unsatisfactory [24]. Since fine-tuning Whisper has been shown to enable transcription of unseen languages [37], we propose to fine-tune Whisper for En-X translation. We fine-tune the audio model in a multi-task style to transcribe English audio and translate it to the other languages. To train Whisper-Flamingo, we freeze the fine-tuned audio model, add the gated cross attention layers and the linear layer on top of the visual features, and train the model on audio-visual inputs.
3 Experiments on LRS3
Table 4: Results for English transcription on LRS3 and En-X Translation on MuAViC. Babble noise is added at 0-SNR (Noisy). One Model= the model translates to all languages with one set of parameters. Test Mod.= inference modalities (Text: T, audio: A, video: V audio-visual: AV). Note that Bilingual AV-HuBERT [12] use a different noise file than us that was not publicly available so the results in noisy conditions are not directly comparable.
Test One Noise En El Es Fr It Pt Ru Avg
[…]
Text-to-Text Translation
Bilingual Transformer [12] T - - 25.8 29.5 27.0 22.6 23.9 17.2 24.3
M2M-100 [38, 12] T - - 24.5 28.7 25.6 21.8 22.2 15.8 23.1
Speech-to-Text Translation (Clean Audio)
Bilingual AV-HuBERT [12] A - - 23.0 27.5 25.1 20.7 20.1 14.7 21.9
Whisper-Small, Fine-tuned A - 2.0 22.4 27.1 24.9 20.9 21.6 15.6 22.1
Whisper-Medium, Fine-tuned A - 2.1 22.9 27.5 26.1 21.9 21.4 15.1 22.5
Whisper-Large, Fine-tuned A - 1.5 23.7 27.9 26.0 21.8 21.4 15.7 22.7
Bilingual AV-HuBERT [12] AV - - 23.4 26.6 25.3 20.7 20.5 14.6 21.9
(Ours) Whisper-Flamingo (Small) AV - 2.0 22.6 27.0 24.7 20.7 21.3 15.5 22.0
(Ours) Whisper-Flamingo (Medium) AV - 1.6 23.8 28.0 26.1 22.5 21.3 16.0 23.0
(Ours) Whisper-Flamingo (Large) AV - 1.3 24.4 27.9 25.9 22.1 21.8 15.7 22.9
Speech-to-Text Translation (Noisy Audio from MuAViC)
Bilingual AV-HuBERT [12] A MuAViC† - 15.9 19.2 17.1 12.9 14.4 10.3 15.0
Whisper-Small, Fine-tuned A MuAViC 17.3 17.5 20.1 19.4 15.3 16.3 11.8 16.7
Whisper-Medium, Fine-tuned A MuAViC 14.8 18.1 22.1 19.8 16.2 17.3 12.1 17.6
Whisper-Large, Fine-tuned A MuAViC 13.8 19.7 23.4 20.4 17.4 17.7 13.3 18.6
Bilingual AV-HuBERT [12] AV MuAViC† - 22.7 24.8 23.8 20.0 20.0 13.7 20.8
(Ours) Whisper-Flamingo (Small) AV MuAViC 10.7 19.0 22.1 21.1 17.1 18.3 13.2 18.5
(Ours) Whisper-Flamingo (Medium) AV MuAViC 8.3 20.7 24.5 21.6 18.8 18.6 13.7 19.6
(Ours) Whisper-Flamingo (Large) AV MuAViC 7.2 21.1 25.4 22.4 19.3 19.9 14.7 20.5
Speech-to-Text Translation (Video only, no audio)
VSP-LLM [39] V - - - 22.7 22.3 17.9 18.7 - -
3.1 Experimental Setup
To train our models, we use LRS3 [10] - the largest, publicly-available AVSR dataset in English (En), sourced from TED talks. We followed AV-HuBERT [4] to create a 433h training set, 1h validation set, and 1h test set. We also combined the LRS3 training videos with 1,326h of English videos from VoxCeleb2 [40] for training. Transcripts for the VoxCeleb2 videos were obtained from Whisper Large-v2 [41]. For En-X translation, we use the MuAViC [12] dataset which has translations of LRS3’s English text into 6 languages: Greek (El), Spanish (Es), French (Fr), Italian (It), Portuguese (Pt), and Russian (Ru).
We use Whisper Small, Medium, and Large-v2 with 244M, 769M, and 1.55B parameters [1]. We extract 80-bin log-Mel spectrograms with a stride of 10ms and window size of 25ms from audio sampled at 16kHz. We extract video features from the AV-HuBERT Large [4] encoder fine-tuned on LRS3 with 325M parameters. For Whisper Large, the gated cross attention layers add 630M parameters, bringing the total number of parameters to 2.5B (including AV-HuBERT). We freeze AV-HuBERT but enable dropout and batch normalization updating during Whisper-Flamingo training. The videos have a frame rate of 25fps and are converted to grayscale. Dlib [42] is used to extract 96x96 crops centered on the lips which are aligned to a reference mean face [43]. During training, a random 88x88 crop is used and the video is flipped horizontally with probability 0.5. For testing, the center 88x88 crop is used.
Table 1 summarizes the hyperparameters for the main experiments. We used A6000 GPUs with 48GB memory. Audio/video samples with similar lengths are batched together, and short samples are 0-padded. AdamW was used as the optimizer  [44]. Following [1], we used SpecAugment [45] (Librispeech-Basic) with Whisper-Large and did not use it with Whisper-Medium. Training was done with PyTorch [46] and PyTorch Lightning [47]. We used the SpecAugment and batch sorter implementations from ESPnet [48].
We train our models in two conditions: without noise (clean) and with noise (noisy). For clean training, we do not add any noise to the audio. For noisy training, we randomly add noise to the audio with a signal-to-noise ratio (SNR) of 0. Following prior work [26, 12], the “natural”, “music” and “babble” noise are sampled from the MUSAN dataset [49], and overlapping “speech” noise is sampled from LRS3 [10]. To select the best checkpoints, we monitor the highest token prediction accuracy on the clean or noisy validation set every 1k steps. Following prior work [12], we use the Fairseq normalizer [50] to remove punctuation and lower-case text before calculating WER. For translation, we use SacreBLEU [51] with the default 13-a tokenizer to calculate BLEU [52].
3.2 Modality Fusion Ablation with Whisper-Medium
We first compared gated cross attention to early and late fusion using Whisper Medium. For early fusion, we duplicate AV-HuBERT’s 25 Hz video features to temporally align them with Whisper’s 50 Hz audio features (after the CNN layers) and use addition to fuse them before Whisper’s Transformer encoder. For late fusion, we use an MLP to fuse the video features with Whisper’s audio features after its Transformer encoder. In both cases, all of Whisper’s parameters are fine-tuned. For audio-only baselines, we use Whisper zero-shot (no fine-tuning) and fine-tuned on LRS3. We test models in both the clean and noisy conditions with babble-noise injected at 0-SNR. The results are shown in Table 2. Fine-tuning audio-only Whisper decreases the noisy WER of the zero-shot model from 22.2% to 12.6%. We then use the fine-tuned model as initialization to train the models with audio-visual fusion. Early-fusion obtained a small improvement in both the clean and noisy WERs. Late-fusion could not fuse the modalities well and performance became worse in both clean and noisy conditions. Finally, Whisper-Flamingo with gated cross attention obtained the best noisy WER, significantly improving the audio-only Whisper fine-tuned baseline from 12.6% to 7.0%, while the clean WER was slightly improved from 1.9% to 1.5%. Freezing Whisper helps retain its strong audio skills while new cross attention layers enable it to integrate the visual modality more effectively.
3.3 Whisper-Flamingo English Speech Recognition
For our main experiments in Table 3, we use Whisper-Large. We report results using beam search with beam size 1 and 15. In noisy conditions, we use babble noise at 0-SNR constructed following AV-HuBERT [26] by adding audio from 30 speakers from LRS3. Results for additional noise types and SNR levels are shown in the Appendix, Section 7.2.
Training on clean audio (without noise). Compared to zero-shot audio-only Whisper-Large, fine-tuning without noise improves the clean ASR WER from 2.1% to 1.0% (using LRS3 433h) and 0.68% (using LRS3+VoxCeleb2 1,759h). Our fine-tuned Whisper-Large achieves SOTA ASR on LRS3 (0.68%) and matches the previous SOTA of 0.7% from Fast Conformer [29]. We then use our fine-tuned Whisper models to initialize audio-visual Whisper-Flamingo and achieve 1.1% / 0.76% AVSR WER using LRS3 433h / LRS3+VoxCeleb2 1,759h respectively. Our audio-visual Whisper-Flamingo achieves SOTA AVSR on LRS3 (0.76%) and matches the current SOTA from Llama-AVSR (0.77%) using LRS3+VoxCeleb2 and Fast Conformer (0.8%) using additional data. Notably, our method only uses 2.5B parameters, while Llama-AVSR uses over 8B parameters. Also, our method achieves better AVSR performance than LP Conformer (0.9%) which was trained on 100k videos. Fine-tuning Whisper on clean audio does not improve the noisy WER, while training Whisper-Flamingo on clean audio improves the noisy WER.
Training on noisy audio. Compared to zero-shot audio-only Whisper-Large, fine-tuning with noise improves the noisy ASR WER from 20.8% to 11.7% (using LRS3 433h) and 11.1% (using LRS3+VoxCeleb2 1,759h). The clean ASR WER is slightly worse compared to Whisper fine-tuned without noise, however, the results are close (0.85% vs 0.68%). We then use our fine-tuned Whisper models to initialize audio-visual Whisper-Flamingo and significantly improve the noisy WER to 5.6% / 5.6% AVSR WER using LRS3 433h / LRS3+VoxCeleb2 1,759h respectively (49.5% relative noisy WER improvement compared to fine-tuned audio-only Whisper). The clean AVSR WER is slightly worse compared to Whisper-Flamingo trained without noise, however, the results are close (0.86% vs 0.76%).
Comparing noisy results to SOTA. Table 3 also shows a comparison with prior audio-visual SSL methods and audio-visual fine-tuning methods on LRS3. Direct comparison in noisy conditions is challenging since different noise datasets were used to generate the babble noise. SSL methods AV-HuBERT [26], u-HuBERT [14], and CMA [27] used LRS3 to generate babble noise, but the noise file they generated was not publicly available. We followed their procedure to generate the noise, so our noisy conditions are similar but not identical. Compared with AV-HuBERT, Whisper-Flamingo achieves better clean performance (0.86% vs 1.4%) and slightly better noisy results (5.6% vs 5.8%), which shows that Whisper-Flamingo is effective at adapting Whisper to the visual features from AV-HuBERT. Moreover, a major advantage of Whisper-Flamingo over AV-HuBERT is improved translation performance (Section 3.4). The best noisy performances are reported by u-HuBERT and CMA; we would like to try them as visual encoders for Whisper-Flamingo, but the weights are not publicly available. Finally, Whisper-Flamingo outperforms other methods in noise which adapt audio-only models through audio-visual fine-tuning [6, 7], including FAVA-USM [6] which was pre-trained on 12M hours of unlabeled audio [2]. However, the babble noise was generated from different datasets making results not strictly comparable.
3.4 Whisper-Flamingo En-X Speech Translation
For these experiments, we fine-tune Whisper and train Whisper-Flamingo with noise. We report results with a beam size of 15.
Audio Results. In Table 4, we show the result of fine-tuning audio-only Whisper-Large for En-X translation using the 6 languages in the MuAViC dataset (“Whisper-Large, Fine-tuned”). Although Whisper was not originally trained for En-X translation, it adapts well to the new task. Testing with clean audio, we achieve an average BLEU score of 22.7, which outperforms the previous SOTA of 21.9 from Bilingual AV-HuBERT. Moreover, our model transcribes En audio (WER of 1.5%) and translates to 6 languages with a single set of parameters, while Bilingual AV-HuBERT fine-tunes separately for each language pair and trains language-specific decoders from scratch. Our model nearly reaches the text-to-text performance from machine translation models using the ground-truth English text; those models achieve average BLEU scores of 23.1 from a multilingual model and 24.3 from bilingual models.
Audio-Visual Results. Once we fine-tune audio-only Whisper for En-X translation, we use it to train Whisper-Flamingo by freezing the weights and adding gated cross attention layers. Testing with clean audio, Whisper-Flamingo slightly outperforms the audio-only model with an average BLEU score of 22.9 and En WER of 1.3%. In noisy conditions, we use multilingual babble noise constructed following MuAViC [12] by adding audio in 9 different languages from 30 speakers. Note that their noise file was not publicly available, so our noisy conditions are similar but not identical. With multilingual babble noise, Whisper-Flamingo significantly outperforms the audio-only Whisper model in average BLEU score (20.5 vs 18.6) and En WER (7.2% vs 13.8%). Compared with the previous SOTA bilingual AV-HuBERT, our audio-only average BLEU is much better (18.6 vs 15.0), but our audio-visual performance is slightly worse (20.5 vs 20.8). However, our models perform both En-X translation and En transcription with a single model, while their models fine-tune separately for each language pair. Finally, we show the results using Whisper-Medium and Whisper-Small: Whisper-Flamingo always does better in noise compared to the audio-only baselines, and performance tends to improve as the model size increases.
Comparison with VSP-LLM. Finally, we compare with VSP-LLM [39], a recent LLM-based approach which uses features from AV-HuBERT as input to an LLM to perform lip-reading and translation using video inputs only (no audio). Even with noise in the audio, our Whisper-Flamingo (Large) outperforms VSP-LLM on all 4 languages evaluated, which shows the advantage of using both audio and video as inputs instead of just video. Also, our model only has 2.5B parameters compared with VSP-LLM’s 7B parameters.
Table 5: Results for English transcription on LRS2. We report results on the original test set (Clean). Hours of unlabeled & labeled audio-visual data used to train each model are shown. Note that we used AV-HuBERT fine-tuned on LRS3 433h.
[…]
Model Unlabeled Labeled ASR AVSR
Audio-Visual methods trained only on transcribed videos
TM-seq2seq [11] - 223 9.7 8.5
CTC/Attention [53] - 380 8.3 7.0
TDNN [54] - 223 6.7 5.9
CM-seq2seq [20] - 380 3.9 3.7
Efficient Conformer [55] - 818 2.4 2.3
AutoAVSR [21] - 3,448 1.5 1.5
Audio-Visual SSL methods
UniAVSR [5] - 380 2.7 2.6
RAVEn [56] 1,759 223 2.5 -
USR [57] 1,759 223 1.9 1.9
Our audio-only Whisper Zero-shot baselines
Whisper-Medium, Beam 1 - - 5.2 -
Whisper-Large, Beam 1 - - 5.5 -
Our models trained with noise
Whisper-Medium FT, Beam 1 - 223 1.3 -
Whisper-Flamingo, Beam 1 1,759 223† - 1.4
4 Experiments on LRS2
We use our framework to conduct experiments on the LRS2 dataset [11]. We followed AutoAVSR [21] to create a 223h training set, 0.6h validation set, and 0.5h test set. We add noise to the audio while training our models. We report the ASR and AVSR WER on the LRS2 test set (clean) in Table 5. Comparing zero-shot Whisper Medium and Whisper Large V2, the medium model performs slightly better (5.2% vs 5.5%), therefore we use Whisper medium for our experiments. Our fine-tuned Whisper and Whisper-Flamingo models have 769M and 1.39B parameters respectively. Compared to zero-shot Whisper, fine-tuned Whisper improves the ASR WER from 5.2% to 1.3%, achieving a new SOTA ASR WER on LRS2. Finally, Whisper-Flamingo achieves SOTA AVSR WER of 1.4%. Whisper-Flamingo outperforms recent SSL method USR [57] (1.9%) trained with the same hours of video and outperforms AutoAVSR [21] (1.5%) trained with more hours of video. Finally, Whisper-Flamingo outperforms Whisper in noisy conditions (shown in the Appendix, Section 7.3).
5 Conclusion
We introduced Whisper-Flamingo, a novel audio-visual model that combines the strengths of AV-HuBERT and Whisper using gated cross attention. Our audio-visual Whisper-Flamingo significantly outperforms audio-only Whisper in noise. We showed that Whisper can be fine-tuned for the new task of X-En translation. Our model performs both En speech recognition and En-X speech translation using one set of parameters while previous methods fine-tune separately on each language. Our method is a generic way of fusing a visual encoder into the decoder of an ASR model to enable AVSR, and it could work with other models trained on more data in the future.
6 Acknowledgments
We thank Alex H. Liu, Mohamed Anwar, and the reviewers for helpful discussion. We thank Videet Mehta for help with the experiments in Table A2. This research was supported by the MIT-IBM Watson AI Lab and an NDSEG Fellowship to A.R.